OpenStudy (gagoside):

what is 4 out of seven times 3 out of 4 this is a fraction btw

OpenStudy (dayakar):

example 3 out of 8 = 3/8

OpenStudy (gagoside):

no its 4/7 times 3/4

OpenStudy (gagoside):

whats the answer plz answer i will fan and medle

OpenStudy (dayakar):

very good times means multiplication

OpenStudy (gagoside):

ya so whats the answer

OpenStudy (dayakar):

try it i will correct

OpenStudy (gagoside):

lol ok

OpenStudy (gagoside):

what is 4/7 multiplied by 3/4

OpenStudy (dayakar):



OpenStudy (gagoside):

ya so what is it

OpenStudy (gagoside):

i think its 16/21

OpenStudy (dayakar):

4 is cancelled both sides then what remains

OpenStudy (gagoside):

3/7 ?

OpenStudy (dayakar):

yes

OpenStudy (gagoside):

mind to help with a few moree??

OpenStudy (gagoside):

i medeled u and faned

OpenStudy (dayakar):

sure

OpenStudy (gagoside):

5*1 3/5

OpenStudy (gagoside):

8?

OpenStudy (dayakar):

is it \[5 \times \frac{ 13 }{ 5 }\]

OpenStudy (dayakar):

\[5 \times 1\frac{ 3 }{ 5 }\]

OpenStudy (gagoside):

second one

OpenStudy (dayakar):

ur answer is correct

OpenStudy (gagoside):

ok how long can u stay and help??

OpenStudy (dayakar):

how many do u have?

OpenStudy (gagoside):

9999999

OpenStudy (gagoside):

byeee

OpenStudy (dayakar):

10^100 sec :)